# Patrick Cheridito and Moritz Weiss

### v1

@@ 6.5 Market with noise & tactical traders
The third and fourth rows of Table 1 and Figure 3(b) display the results for the market consisting of noise and tactical traders. Both demonstrate the LN algorithm’s superior performance relative to the benchmark algorithms for small and large position sizes. The performance of the LN algorithm is slightly better than the DR algorithm. As before, for 20 lots, we observe bumps in the histogram at one and two ticks above the best bid price for the LN algorithm. For $60$ lots, the LN algorithm’s histogram has less left skew than the benchmark densities, indicating that it cuts losses when prices move against it. The presence of tactical traders reacting to order book imbalances introduces indirect market impact. Consequently, posting the entire position passively at the beginning, as done by the SL algorithm, triggers the tactical traders to trade in the other direction, leading to the worst performance. Spreading the order out over the whole time horizon, as done by the TWAP algorithm, is better. Still, the LN algorithm performs best, suggesting intelligent placement strategies.
@@ 6.6 Market with noise & tactical & strategic traders
The fifth and sixth rows of Table 1 and Figure 3(c) show the results for the market consisting of noise, tactical, and strategic traders and demonstrate the strong performance of the LN algorithm. The presence of a strategic trader introduces an upward or downward drift in the market. The LN algorithm detects this drift using the historical market and limit order flow features and the mid-price drift feature defined in Section 3.1. The performance of the LN algorithm is slightly better than the DR algorithm. The relative shift of the LN algorithm’s histogram compared to those of the benchmark algorithms suggests that it trades more aggressively when the strategic trader is selling and more patiently when the strategic trader is buying.
@@ 7 Conclusion
We have developed an execution algorithm based on a novel actor-critic policy gradient approach that uses the logistic-normal distribution to model random allocations. Its state space captures all relevant market features, including current and past states of the order book and queue positions. The action space allows the algorithm to place market and limit orders across multiple price levels, adapting to changing market conditions. The algorithm outperforms competitive benchmarks in realistic simulations that account for direct and indirect market impact. This is the first time the logistic-normal distribution has been combined with the actor-critic algorithm. Our results suggest that this is a better approach than using the Dirichlet distribution since our algorithm shows better results and faster convergence speeds.
The strength of our results is inherently tied to the realism of the market simulation. While our simulations capture key stylized facts of limit order book markets, they do not fully replicate real-world conditions. Nevertheless, our method remains broadly applicable and can be used with any market simulation framework and institutional investors may combine our method with their own proprietary market simulations.
Appendix A Parameters and stationary order book shapes in the market simulation
@@ A.1 Parameters for market simulations
We assume that the noise and tactical traders submit orders (market orders, limit orders and cancellations) up to level $D=30,$ and that all of their orders have the same shifted half-normal distribution with parameters $\sigma^{M}=\sigma^{L,k}=\sigma^{C,k}=2,$ for $k\in\{1,2,\dots,D\}.$ We cap large order sizes at five standard deviations because very large order sizes are unrealistic. We always start the simulations at $-\Delta t=-15s,$ from an equilibrium state and run them until time $T=150s.$ The equilibrium states are explained in more detail in Appendix A.2 below. The LN-RL, DR-RL, and all benchmark algorithms trade during the time interval [0, T], updating their decisions at the times $n\Delta t,n=0,1,…,N,$ for $\Delta t=15s$ and $N=10.$ The parameters are summarized in Table 2.
@@
start time for noise, tactical & strategic traders $-\Delta t$ $-15s$
@@
$N$ 10
$\sigma^{M}=\sigma^{L,k}=\sigma^{C,k}$ 2
@@
The intensities of limit order arrivals $\lambda^{L,k}$ and cancellations $\lambda^{C,k}$ for the noise traders are given in Table 3. The market order intensity is $\lambda^{M}=0.1237.$ Those intensities are taken from the paper Abergel and Jedidi (2013) with one modification. Instead of the cancellation intensities in that paper, we increase the cancellation intensities by a factor of $100$ This leads to smaller queue sizes, which makes the simulation more efficient. At $-\Delta t$, we set the limit order book equal to its equilibrium shape. It is displayed in Figure 5(a). The average total traded volume in this market is 100 lots.
@@
For the tactical traders, we choose a damping factor $c=0.65$ and constant imbalance reaction factors $d^{L,k}=d^{C,k}=d^{M}=2$ for $k\in\{1,2,\dots D\}.$ To compensate for the presence of tactical traders, we reduce the intensities of the noise traders by 15%. With this intensity reduction, the average total traded volume over the trading horizon is roughly the same for the market with noise traders only and the market with noise and tactical traders. We start the simulation at $-\Delta t$ from its equilibrium state, displayed in Figure 5(b).
@@
We combine noise, tactical, and strategic traders for the third market. The parameters for noise and tactical traders remain the same as in the previous market simulations, except that we reduce the intensities of the noise and tactical traders by 10% to account for the presence of the strategic trader. This way, the total average traded volume is roughly the same for all market environments. The strategic trader sends a limit order of size $\nu^{L}=2$ every $\Delta t^{L}=3$ seconds and a market order of size $\nu^{M}=1$ every $\Delta t^{M}=3$ seconds. Due to the presence of the strategic trader, this market does not have an equilibrium state. At time $-\Delta t,$ we start the order book from the same shape as in the market with noise and tactical traders. The trade direction of the strategic trader (buying or selling) is drawn randomly with probability $1/2$ at the time $-\Delta t.$
Small and large initial positions
The total average traded volume in all three simulated markets is roughly the same. Therefore, the results for the initial positions of 20 and 60 lots in Table 1 are comparable across all markets, as in all markets this accounts for roughly 20% and 60% of the total volume.
@@ A.2 Stationary order book shapes
We always start the market simulation from an equilibrium state of the limit order book. The equilibrium states are described by the average queue sizes and the average spread:
@@
Here, for $k\in\{1,\dots,D\}$ the quantity $\tilde{v}^{b,k}$ is the average volume $k-1$ ticks below the best bid price and $\tilde{v}^{a,k}$ is the average volume $k-1$ ticks above the best ask price. Since the average spread $\tilde{s}$ in our simulations is always close to one tick, we start all simulations with an initial bid price $p^{b}(-\Delta t)=1000$ and an initial ask price $p^{a}(-\Delta t)=1001$. An asset with an average spread close to one tick is called large tick asset.
@@
Appendix B Hyperparameters and implementation details for the actor-critic algorithm
We implemented an actor-critic reinforcement learning algorithm that solves a dynamic allocation problem using the logistic-normal distribution. Our implementation is based on the implementation of the proximal policy optimization algorithm (PPO) by Huang et al. (2022); see https://github.com/vwxyzjn/cleanrl. We implemented the following changes. First, we are not using clipping. Second, we implemented a custom logistic-normal action policy. Third, we implemented the custom initialization of the mean and covariance matrix and the scaling of the covariance matrix during the training loop described in Section 4.3.
In this section, we provide a detailed description of the algorithm’s hyperparameters. They are summarized in Table 4 and Table 5.
@@ Appendix B Hyperparameters and implementation details for the actor-critic algorithm
Number of hidden layers $2$
@@
Nodes of output layer for value function network $1$
@@
For the value function $V_{\vartheta}$ and the policy $\pi_{\theta},$ we use two separate neural networks with two hidden layers with $128$ nodes. The hidden layers have a $\tanh$ activation function and the last layers have no activation function. The last layer of the policy network has $K$ nodes corresponding to the mean vector $\mu=(\mu^{0},\dots,\mu^{K-1})$ of the normal distribution with density $\varphi_{\theta}.$ The last layer of the value function network has one node. We initialize the weights of both neural networks with an orthogonal initialization scheme; see e.g. Saxe et al. (2013). For the last layer of the policy network, we set the bias to $b=(-1,-1,\dots,-1)\in\mathbb{R}^{K}$ and initialize the weights using orthogonal initialization with a gain factor $10^{-5}.$ Therefore, the weights of the last layer of the policy network are close to zero and the output is effectively determined by the bias term. For all other layers in the policy and value function neural networks, we use an orthogonal initialization with a gain factor $0.01$ to initialize the weights and set the bias terms to zero.
@@
[…]
@@
We use the Adam optimizer (Kingma and Ba, 2017) with a learning rate (also referred to as step-size) of $\eta=0.0005.$ In all experiments, we make $H=400$ gradient steps. We set the initial covariance matrix $\Sigma=(\sigma^{ij})_{i,j\in\{0,\dots,K-1\}}\in\operatorname{\mathbb{R}}^{K\times K}$ of the normal distribution $\varphi_{\theta}$ equal to the unit matrix in $\operatorname{\mathbb{R}}^{K\times K}.$ We choose the initial and final variance $\sigma^{\text{init}}=1.0$ and $\sigma^{\text{final}}=0.1$ and, in each training iteration $i\in\{1,\dots,H\},$ reduce the diagonal entries according to the schedule (12).
Appendix C Implementation details for Dirichlet reinforcement learning algorithm
@@
Let $\alpha=(\alpha^{0},\alpha^{1},\alpha^{2},\dots,\alpha^{K})$ with $\alpha^{k}>0$. The Dirichlet distribution $\text{Dir}(\alpha)$ is defined on the probability simplex:
@@
The probability density function is defined by
$\pi(a^{0},a^{1},\dots,a^{K})=\frac{1}{B(\alpha)}\prod_{k=0}^{K}(a^{k})^{\alpha^{k}-1}\;\text{ with }\;B(\alpha)=\frac{\prod_{k=0}^{K}\Gamma(\alpha^{k})}{\Gamma\left(\sum_{k=0}^{K}\alpha^{k}\right)}.$
@@
As for the logistic-normal distribution, we use the policy gradient theorem (4) to train the algorithm. The parameters of the density $\alpha^{0},\alpha^{1},\dots,\alpha^{K}$ are the output of a neural network. To ensure that the parameters are positive, we use the soft-plus function, defined as $\exp(1+\log(\cdot)),$ in the output layer as an activation function. We choose the soft-plus function over the exponential function to prevent the parameters from becoming excessively large, which could lead to a concentrated policy. Furthermore, when initializing the policy neural network, we choose the weights of the output layer close to $(1,1\dots,1,10).$ This ensures that the action $a^{K}$ has more weight than the other actions. Therefore, at the beginning of the training cycle, the algorithm will not place many orders, observing full trajectories. The neural network architecture for the value function network is the same as above.
Appendix D Feature normalization
The features described in Section 3.1 are normalized in our implementation. We generally aimed to normalize all features to the same scale by transforming them to the interval $[-1,1]$ or $[0,1].$ We present each feature’s normalized counterpart in the following bullet points.

### v2

@@ 6.6 Market with noise & tactical traders
The third and fourth rows of Table 1 and Figure 4(b) show the results for the market with noise and tactical traders. Both demonstrate the LN algorithm’s superior performance relative to the heuristic benchmark algorithms for small and large position sizes. The LN algorithm performs better than the DR algorithm for an initial position of 20 lots and slightly worse for 60 lots. As before, for 20 lots, we observe bumps in the histogram one and two ticks above the best bid price for the LN algorithm. The presence of tactical traders reacting to order book imbalances introduces indirect market impact. Consequently, posting the entire position passively at the beginning, as the SL algorithm does, triggers tactical traders to place sell orders, leading to the worst performance for both initial positions. Spreading the order out over the entire execution interval, as done by the TWAP algorithm, works better. Still, the LN algorithm performs better, suggesting intelligent order placement strategies that don’t provoke a reaction from tactical traders.
@@ 6.7 Market with noise & tactical & strategic traders
The fifth and sixth rows of Table 1 and Figure 4(c) show results for the market consisting of noise, tactical, and strategic traders and demonstrate a strong performance of the LN algorithm. The presence of the strategic trader introduces an upward or downward drift in the market. The LN algorithm detects this drift using the historical market, limit, and cancellation order flow features and the mid-price drift feature defined in Section 3.1. The LN algorithm outperforms the DR algorithm, but the DR algorithm’s returns have a lower standard deviation when the initial position is 60 lots. The LN algorithm has less skew to the left than the DR algorithm for an initial position of 20 lots, and more skew to the right for an initial position of 60 lots. The relative shift of the LN algorithm’s histogram compared to those of the heuristic benchmark algorithms suggests that it trades more aggressively when the strategic trader is selling and more patiently when the strategic trader is buying.
6.8 Analyzing the algorithms’ trading behavior
In this section, we run two experiments that illustrate the trading behavior of the LN and DR algorithms. Averages and standard deviations are computed over 10,000 simulations, after the algorithms have been trained. In addition to these experiments, we also conduct an ablation study in Appendix C.4 to assess feature importance.
The first experiment concerns the algorithms’ inventories. Figure 6 shows the average inventory curves for the LN and DR algorithms in bold. The inventory is the number of lots $M(t)$ that the algorithm still has to sell at time $t,$ defined in Section 3. The green curves correspond to the LN algorithm, whereas the red curves correspond to the DR algorithm. The shaded areas indicate standard deviations. We observe that the LN algorithm trades 60 lots more patiently in the market with both noise and tactical traders than in the market with noise traders only. In contrast, the DR algorithm exhibits roughly the same trading speed in both markets. In the market with noise, tactical, and strategic traders, we observe significantly larger inventory standard deviations for both algorithms. This is because the algorithms liquidate faster or slower depending on whether the strategic trader is buying or selling. For the 60-lot case in the market with noise, tactical, and strategic traders, we observe that the LN algorithm holds back more inventory at the beginning of the execution period than the DR algorithm, indicating that the algorithms have different order placement strategies.
(a) Noise Traders
(b) Noise & Tactical Traders
(c) Noise, Tactical, & Strategic Traders
Figure 6: Average inventory curves are displayed in bold, and the shaded areas around them indicate standard deviations. The green curves correspond to the LN algorithm, whereas the red curves correspond to the DR algorithm. Deeper color tones correspond to larger initial positions.
The second experiment concerns the algorithms’ actions. Figure 7 shows the average actions at the initial time $t=0$ for the LN and DR algorithms, and for different initial positions. Figure 7(a) demonstrates that both algorithms place limit orders aggressively, using actions $a_{1}$ and $a_{2}$ in the market with noise traders only. This is because limit orders placed at the best ask price have little market impact, given the absence of tactical traders. Interestingly, for 60 lots, they also send market orders using action $a_{0}.$ Figure 7(b) demonstrates a more balanced passive placement in the market with noise and tactical traders, with greater use of actions $a_{5}$ and $a_{6}$ for both the LN and DR algorithms. This occurs because placing large limit orders near the best ask price triggers reactions from tactical traders. For 60 lots, both algorithms allocate many orders to action $a_{6}$ to avoid provoking such reactions. Figure 7(c) shows large allocations to both market and defensive limit orders, using actions $a_{0},a_{5}$ and $a_{6},$ in the market with noise, tactical, and strategic traders. This is because it is more profitable to send a market order when the strategic trader is selling, as the price is more likely to move down. When the strategic trader is buying, it is better to place a patient limit order far from the best ask using action $a_{5}$ or to withhold orders from the market using action $a_{6}$ to wait for the price to rise.
(a) Noise Traders
(b) Noise & Tactical Traders
(c) Noise, Tactical, & Strategic Traders
Figure 7: Average actions at time $t=0$ for the LN and DR algorithms across all simulated markets. The green and red bars correspond to the average actions chosen by the LN and DR algorithms, respectively. The upper panel shows the actions for 20 lots, whereas the lower panel shows them for 60 lots.
@@ 7 Conclusion
In this paper, we develop a trade execution algorithm based on a novel actor-critic policy gradient approach that uses the logistic-normal distribution to model random allocations. Its state space captures all relevant market features, including current and past states of the order book and queue positions. The action space allows the algorithm to place market orders and limit orders across multiple price levels, adapting to changing market conditions. The algorithm outperforms competitive benchmarks in realistic simulations that account for direct and indirect market impact. This is the first time the logistic-normal distribution has been combined with the actor-critic algorithm. Our simulations capture key stylized facts of limit order book dynamics. But our method can be used with any market simulation framework.
Appendix A Implementation Details
@@ A.1 Parameters for market simulations
We assume that the noise and tactical traders submit orders (market orders, limit orders and cancellations) up to level $D=30,$ and that all of their orders have the same shifted half-normal distribution with parameters $\delta^{M}_{\text{noise}}=\delta^{L,k}_{\text{noise}}=\delta^{C,k}_{\text{noise}}=2,$ and $\delta^{M}_{\text{tactical}}=\delta^{L,k}_{\text{tactical}}=\delta^{C,k}_{\text{tactical}}=2,$ for $k\in\{1,2,\dots,D\}.$ We cap order sizes at five standard deviations. We always start the simulations at $\Delta t=-15s,$ from an equilibrium state and run them until time $T=150s.$ The equilibrium states are explained in more detail in Appendix A.2 below. The LN algorithm, DR algorithm, and the two heuristic benchmark algorithms trade during the time interval [0, T], updating their decisions at the times $n\Delta t$ for $n=0,1,\dots,N-1,$ with $\Delta t=15s$ and $N=10.$ The simulation parameters are summarized in Table 2.
@@
start time for noise, tactical & strategic traders $\Delta t$ 15s-15s
@@
NN 10
$\delta^{M}_{\text{noise}}=\delta^{L,k}_{\text{noise}}=\delta^{C,k}_{\text{noise}}$ 2
$\delta^{M}_{\text{tactical}}=\delta^{L,k}_{\text{tactical}}=\delta^{C,k}_{\text{tactical}}$ 2
@@
The intensities of limit order arrivals $\lambda^{L,k}$ and cancellations $\lambda^{C,k}$ for the noise traders are given in Table 3. The market order intensity is $\lambda^{M}=0.1237.$ Those intensities are taken from the paper Abergel and Jedidi (2013) with one modification. Instead of the cancellation intensities in that paper, we increase the cancellation intensities by a factor of $100.$ This leads to smaller queue sizes, which makes the simulation more efficient. At $\Delta t$, we set the limit order book equal to its equilibrium shape. It is displayed in Figure 8(a).
@@
For the tactical traders, we choose a damping factor $c=0.65$ and constant imbalance reaction factors $d^{L,k}=d^{C,k}=d^{M}=2$ for $k\in\{1,2,\dots D\}.$ To compensate for the presence of tactical traders, we reduce the intensities of the noise traders by 15%. With this intensity reduction, the average total traded volume over the trading horizon is roughly the same for the market with noise traders only and the market with noise and tactical traders. We start the simulation at $\Delta t$ from its equilibrium state, displayed in Figure 8(b).
@@
We combine noise, tactical, and strategic traders for the third market. The parameters for noise and tactical traders remain the same as in the previous market simulation. The strategic trader sends a limit order of size $\nu^{L}=2$ every $\Delta t^{L}=3$ seconds and a market order of size $\nu^{M}=1$ every $\Delta t^{M}=3$ seconds. Due to the presence of the strategic trader, this market does not have an equilibrium state. At time $\Delta t,$ we start the order book from the same shape as in the market with noise and tactical traders. The trade direction of the strategic trader (buying or selling) is drawn randomly with probability $1/2$ at the time $\Delta t.$
@@ A.2 Stationary order book shapes
We always start the market simulation from an equilibrium state of the limit order book. The equilibrium states are described by the average volumes and the average spread:
@@
Here, for $k\in\{1,\dots,D\}$ the quantity $\tilde{v}^{b,k}$ is the average volume k1k-1 ticks below the best bid price and $\tilde{v}^{a,k}$ is the average volume k1k-1 ticks above the best ask price. Since the average spread $\tilde{s}$ in our simulations is always close to one tick, we start all simulations with an initial bid price $p^{b}(-\Delta t)=1000$ and an initial ask price $p^{a}(-\Delta t)=1001$.
@@
A.3 Traded volumes
In this section, we briefly document traded volumes and event statistics, which are computed over 10,000 trading simulations for all markets with a time horizon of 150 seconds. The statistics are documented in Table 4, which consists of five columns. The second column shows the average number of events. An event is a market order placement, a limit order insertion, or a cancellation. The third column shows the average traded volume. The fourth and fifth columns show the average buy and sell volumes, which sum to the total volume. There are many more events than traded volume, because limit order insertions and cancellations do not increase the traded volume. The total average traded volume in the first two markets is roughly the same. The presence of the strategic trader increases the average traded volume in the market with noise, tactical, and strategic traders.
Table 4: Average number of events (limit, market, and cancellation orders), the average number of trades, and the average total buy and sell volume for all market simulations.
market #events total volume buy volume sell volume
noise 1162 95 48 47
noise & tactical 1158 98 49 49
noise & tactical & strategic 1389 149 73 76
Appendix B Implementation details for actor-critic reinforcement learning algorithms
B.1 Logistic-normal algorithm
We implement an actor-critic reinforcement learning algorithm combined with the logistic-normal distribution for the trade execution problem. Our implementation is based on the implementation of the proximal policy optimization algorithm (PPO) by Huang et al. (2022); see https://github.com/vwxyzjn/cleanrl. We implement the following changes. First, we are not using clipping. Second, we implemented a custom logistic-normal action policy. Third, we implement a custom initialization of the mean of the underlying normal distribution and variance scaling as in 13. The algorithm’s hyperparamters are summarized in Table 5 and Table 6.
@@ B.1 Logistic-normal algorithm
Number of hidden layers 22
@@
Nodes of output layer for value function network 11
@@
For the value function $V_{\vartheta}$ and the policy $\pi_{\theta},$ we use two separate neural networks with two hidden layers with $128$ nodes. The hidden layers have a $\tanh$-activation function and the last layers have no activation function. The last layer of the policy network has $K$ nodes corresponding to the mean vector $\mu=(\mu^{0},\dots,\mu^{K-1})$ of the normal distribution with density $\varphi_{\theta}.$ The last layer of the value function network has one node. We initialize the weights of both neural networks with an orthogonal initialization scheme; see e.g. Saxe et al. (2013). For the last layer of the policy network, we set the bias to $b=(-1,-1,\dots,-1)\in\mathbb{R}^{K}$ and initialize the other weights using orthogonal initialization with a gain factor $10^{-5}.$ Therefore, the weights of the last layer of the policy network are close to zero, except for the bias term, and the output is effectively determined by the bias term. With our choice of the bias term, the action $a^{K}$ is more likely than the other actions at initialization, as explained in (18). For all other layers in the policy and value function neural networks, we initialize the bias terms to zero and use an orthogonal initialization scheme with a gain factor $0.01$ for all other weights.
@@
We use a vectorized architecture to collect samples from the market simulation. We use $128$ parallel markets, using 128 CPUs, where each market collects 10 trajectories. This results in a total number of $\tau=\text{1,280}$ trajectories. In each training iteration, we sample a batch of trajectories
@@
We use the Adam optimizer (Kingma and Ba, 2017) with a learning rate (also referred to as step-size) of $\eta=0.0005.$ In all experiments, we make $H=400$ gradient steps. We choose the initial and final variance $\sigma_{\text{init}}=1.0$ and $\sigma_{\text{final}}=0.1,$ and reduce action variance following the schedule (13).
Computational resources
The experiments were conducted on a workstation equipped with an AMD EPYC 7763 64-Core Processor with 128 threads, and a single NVIDIA GeForce RTX 4090 GPU with 24 GB VRAM. The system contains four NUMA nodes and 256 MiB of L3 cache, running with CUDA version 12.8 and driver 570.144.
Training times
Simulating the order book is an expensive operation. We need to keep track of all orders on all price levels. Whenever new market, limit, or cancellation orders are received, they must be matched with the resting limit orders in the order book. To achieve this, we employ dictionaries to keep track of resting limit orders. Market orders require explicit iteration over price levels, whereas cancellations and limit orders can be processed without for-loops. Due to these computational complexities, training times for all our algorithms are relatively long. They are summarized in Table 7 for the LN and DR algorithms. Training times are measured in hours and minutes. We notice that the DR algorithm trains faster than the LN algorithm in most cases. However, training duration also depends on the overall load on our computing machines, so these measurements should be interpreted only as rough indicators of computational cost. The DR algorithm was likely trained overnight, when system load was lower.
Table 7: Training times for the LN and the DR algorithms in hours and minutes ($h$:$m$). The results are displayed for the different market environments and initial positions.
Market Lots DR LN
Noise 20 1:35 1:28
60 1:13 1:29
Noise & Tactical 20 1:36 1:28
60 1:19 1:43
Noise & Tactical 20 1:32 1:52
& Strategic 60 1:26 1:55
B.2 Dirichlet algorithm
@@
Let $\alpha=(\alpha^{0},\alpha^{1},\alpha^{2},\dots,\alpha^{K})$ with $\alpha^{k}>0,$ for $k=0,1,\dots,K.$ The Dirichlet distribution $\text{Dir}(\alpha)$ is defined on the probability simplex:
@@
Setting $\bar{\alpha}=\sum_{k=0}^{K}\alpha^{k},$ the probability density function is defined by
$\pi(a^{0},a^{1},\dots,a^{K})=\frac{1}{B(\alpha)}\prod_{k=0}^{K}(a^{k})^{\alpha^{k}-1}\;\text{ with }\;B(\alpha)=\frac{\prod_{k=0}^{K}\Gamma(\alpha^{k})}{\Gamma\left(\bar{\alpha}\right)}.$
Let $a$ be a random variable with Dirichlet distribution, then the expected value for each component is given by
$\operatorname{\mathbb{E}}\big[a^{k}\big]$ $=\frac{\alpha^{k}}{\bar{\alpha}},\quad\text{for }k=0,\dots,K.$ (21)
The covariance for $i,j\in\{0,1,\dots,K\}$ is given by
$\mathrm{Cov}\left[a^{i},a^{j}\right]=\begin{cases}-\frac{\alpha^{i}\alpha^{j}}{{\bar{\alpha}}^{2}({\bar{\alpha}}+1)}\quad&i\neq j,\\ \frac{\alpha^{i}(\bar{\alpha}-\alpha^{i})}{{\bar{\alpha}}^{2}(\bar{\alpha}+1)}\quad&i=j.\end{cases}$ (22)
@@
As for the logistic-normal distribution, we use the policy gradient formula (4) to train the algorithm. The neural network architecture for the value function network is identical to that of the LN algorithm, which was defined in Section B.1. The parameters of the density $\alpha_{\theta}=(\alpha^{0}_{\theta},\alpha^{1}_{\theta},\dots,\alpha^{K}_{\theta})$ are the output of a neural network (policy network) with weights $\theta$. The network architecture is the same as for the LN algorithm, except for the output layer. To ensure that the outputs are positive, we use the soft-plus function, defined as $\mathrm{softplus}(x)=\log(1+\exp(x)),$ as an activation function in the output layer. We choose the soft-plus function over the exponential function to prevent the parameters from becoming excessively large, which could lead to a concentrated policy due to (22). As for the logistic-normal distribution, we initialize the weights of the final layer of the policy network with an orthogonal initialization scheme using a gain factor of $10^{-5},$ leading to small weights. We set the bias of the final layer to $b=\textrm{softplus}^{-1}(1,1,\dots,1,10)\in\operatorname{\mathbb{R}}^{K+1},$ where the inverse is applied to each component. Thus, after applying the soft-plus activation function, the output is close to $\alpha_{\theta}=(1,1,\dots,1,10)\in\mathbb{R}^{K+1}.$ Then (21) implies
$\mathbb{E}[a_{\theta}^{K}]\approx 10\mathbb{E}[a_{\theta}^{k}]\quad\text{for }k=0,\dots,K-1.$
Therefore, the action $a_{\theta}^{K}$ (not placing orders) is more likely than the other actions $a_{\theta}^{0},\dots,a_{\theta}^{K-1}.$ Hence, the algorithm is more likely to observe full trajectories up to terminal time $T$ at the beginning of the training cycle, which helps to prevent it from getting stuck in local minima. When evaluating the algorithm, we use a deterministic version. More precisely, we use the expected value (21) with parameters $\alpha_{\theta}$ as actions, where $\theta$ are the final trained weights of the neural network.
B.3 Feature normalization
The features described in Section 3.1 are normalized in our implementation. We aim to normalize all features to a common scale by transforming them to the interval [1,1][-1,1] or $[0,1]$. We present the normalized counterpart of each feature in the following bullet points. Furthermore, we introduce an additional feature.